What was found
AT TIME 14:13-Z ON BOTH HYD PUMPS WITH RES LOW PRESS LT ON. AUX PUMP WOULD ONY BUILD UP TO 1000 PSI, QUANITY STAYED AT 3/4. FLAPS SHOWED A SPLIT WHEN SYSTEM PRESSURIZED IN THE BYPASS POSS. FOUND BLOWN O-RING ON SLOT CONTROL VALVE. REPLACED O-RING ON SLOT CONTROL VALVE SERVICED HYD SYS AS REQUIRED PERFORMED ENG RUN ON NR 2 AND NR 3 ENG FOR OPS CK HYD PUMPS, PERFORMED FLT CONTROL CKS AND LEAKS CKS OF NEW O-RING, OPS CK GOOD ON ENG DRIVEN PUMPS, FLT CONTROL LEAKS CK GOOD. NO LEAKS OR DEFECTS NOTED.